OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 28 feet, what is its lateral surface area? Use π = Answer 176 ft2 392 ft2 686 ft2 1232 ft2

OpenStudy (lgbasallote):

LSA = 2 pi rh A = pi r^2 r = sqrt (A/pi) r = sqrt (154/pi) LSA = 2 pi (sqrt 154/pi)(28) do you know the answer already? since you closed it...

OpenStudy (anonymous):

no i dont know it i didnt mean to

OpenStudy (lgbasallote):

well then..just input those values in the calculator now..^_^

OpenStudy (anonymous):

is it 1232?

OpenStudy (lgbasallote):

yeah

OpenStudy (lgbasallote):

you got it ^_^